# Periodic orbits for square and rectangular billiards

### v1

@@ Abstract
keywords:
periodic orbit; mathematical billiard; Euler’s totient function.
articletype: RESEARCH ARTICLE
@@ 1 Introduction
Mathematical billiards shares many similarities to the game of billiards in reality. Fundamentally, both comprise a billiard ball moving in a straight line until it strikes a side of the table. Mathematical billiards ignores friction, takes the billiard ball as a point mass and assumes purely elastic collisions. As a consequence, the billiard trajectory will satisfy the law of reflection (the angle of incidence equals the angle of reflection) when it collides with a boundary. Therefore, any billiard trajectory is uniquely determined by its initial position and direction of motion. A real billiard table is almost always rectangular, but mathematical billiard tables can have arbitrary shapes and dimensions. First posed as Alhazen’s problem in optics by Ptolemy in 150 AD (Baker 1881), mathematical billiards is much more than a fun and interesting game; it turns out that the mere shape of the table distinguishes mathematical billiards into three different classes—elliptic, hyperbolic, and parabolic—which are well known as classes in different fields of mathematics, not the least in dynamical systems theory (Sinai 1991; Katok 2009). Mathematical billiards has extensively been studied from the perspective of ergodic theory in dynamical systems, as well as algebraic geometry (moduli spaces) and Teichmüller theory; for example, see the textbooks by Tabachnikov (2005), Chernov and Markarian (2006), or Rozikov (2018). Nevertheless, a myriad of unsolved open problems in mathematical billiards remain (Dettmann 2011; Gutkin 2012; Bialy et al. 2022; Kaloshin and Sorrentino 2022). Indeed, there are many fundamental theory questions in mathematics that can be recast as a billiard problem (Sinai 2000, 2004), as well as a plethora of applications, including but not limited to the computation of $\pi$ (Galperin 2003), mechanics (Arnold 1989), quantum computing (Fredkin and Toffoli 1982), pouring problems (Polster and Ross 2012), Benford’s Law (Tabachnikov 2005), diffusion in the Lorentz Gas (Dettmann 2014) and the Riemann hypothesis (Bunimovich and Dettmann 2005).
@@
Obviously, singular orbits only exist on tables with vertices and they are often ignored or classified as non-periodic (Boshernitzan et al. 1998). If a billiard trajectory is periodic, its period is given by the total number of unique collisions. A natural question to ask is whether periodic orbits always exist or whether this depends on the shape of the billiard table. Furthermore, what periods are possible? In fact, the very question whether periodic orbits exist on any polygon is still an open problem! It is listed as Problem 3 in Katok’s five most resistant problems in dynamics (Katok 2004). Even the case of a triangular billiard table is considered impenetrable, although the Triangular Billiards Conjecture, which states that every triangular table has a periodic billiard path, is widely believed to be true (Schwartz 2009).
The question about which periods are possible has been treated for specific polygonal billiard tables. In particular, Baxter and Umble (2008) classified all possible periodic orbits for an equilateral triangle. In this paper, we consider square or rectangular billiard tables and we similarly give a complete classification of which periods and types of periodic orbits are possible for a square or rectangle. The square is perhaps the simplest piecewise-smooth billiard table, which means that it is introduced early on in billiard textbooks, typically alongside the smooth, circular billiard; for example, see (Tabachnikov 2005; Chernov and Markarian 2006; Rozikov 2018). As a consequence, results for the square billiard tend to be skimmed over, despite its behaviour being distinctly different from circular and other smooth, convex tables. Existence of periodic orbits on the square and rectangular billiard is easily verified by explicit examples. Figure 1 shows three examples of periodic orbits on the square. Here, the square table is represented by the (blue) boundary with vertices A, B, C and D, the periodic orbits are indicated as directed (black) lines that collide with the boundary at points identified by thick (red) dots. By counting these points of collision, we see that the periods are two, four and six in the respective panels (a), (b) and (c). Note that the period-two orbit in panel (a) is readily produced on a rectangle as well; for the period-four and -six orbits this is less obvious.
@@
Interestingly, none of the textbooks show that the square and rectangular billiard tables do not admit periodic orbits with odd periods. This can be proven using cutting sequences, as done by Davis (2016), but we provide a different proof that harmonises well with the periodicity classification presented in this paper. Furthermore, we also discuss how to construct periodic orbits with prespecified periods and connect our classification of periodic orbits with Euler’s totient function from number theory (Loomis et al. 2008).
This paper is organised as follows. In the next section, we define the different classes of periodic orbits and clarify what we mean when two periodic orbits with the same period are different. Here, we also explain the powerful technique of unfolding (Fox and Kershner 1936; Zemlyakov and Katok 1975) that is used in the classical proof of the existence of periodic orbits on the square or rectangular billiard table. We present alternative and intuitive proofs for periodicity and non-existence of odd periodic orbits in Section 3. In Section 3.2, we count and fully classify all possible periodic orbits. We treat singular orbits in Section 3.3, where we show how different types of singular orbits form boundaries between different families of periodic orbits. In Section 4, we extend the results for the square to the rectangular billiard table. We conclude in Section 5 with a discussion of future work.
@@ 2 Setting and definitions
Proof.
Since $\alpha_{0}\in(0,\frac{\pi}{2}]$, the outgoing line from $P_{0}$ on the side AB will either end on the adjacent side BC or on the opposite side CD. The sides AB and CD are parallel, because ABCD is either a square or a rectangle. Hence, if the outgoing line from $P_{0}$ ends on CD, is does so at the same angle $\alpha_{0}$; see Figure 1(a) and 1(c) for examples. If the outgoing line from $P_{0}$ ends on the adjacent side BC then it does so at an angle $\alpha_{1}$ such that $\alpha_{0}+\frac{\pi}{2}+\alpha_{1}=\pi$, because the outgoing line forms a right triangle with the sides AB and BC; see Figure 1(b) and 1(c) for examples. Therefore, we have $\alpha_{1}=\frac{\pi}{2}-\alpha_{0}$ as claimed. The end point of the outgoing line from $P_{0}$ is another collision point $P_{1}$, unless it is a vertex, in which case the billiard trajectory terminates and the result holds. By rotating the table and relabelling the vertices, we can consider the billiard trajectory as being generated by the pair $\langle P_{1},\alpha_{1}\rangle$, with $\alpha_{1}=\alpha_{0}$ or $\alpha_{1}=\frac{\pi}{2}-\alpha_{0}$. Using the same arguments, we find that a collision with the (original) side DA also occurs at angle $\frac{\pi}{2}-\alpha_{0}$. It follows that all collisions with sides AB and CD occur at angle $\alpha_{0}$ and all collisions with sides BC and DA occur at angle $\frac{\pi}{2}-\alpha_{0}$. ∎
@@
Definition 2.2 (Equivalence class $\mathcal{C}_{K}(p)$ of period-$K$ orbits).
@@ 2.1 Unfolding of billiard trajectories
Following a billiard trajectory on a table ABCD can be difficult if there are many collisions, because of a myriad of intersecting lines. The technique of unfolding transforms the billiard trajectory: rather than reflecting collisions with the sides, the entire table reflects, so that the billiard trajectory remains a straight line. For so-called rational, planar, polygonal billiard tables, which have polygonal angles that are rational multiples of $\pi$, this approach relates mathematical billards to the theory of geodesic flow on Riemann surfaces (Fox and Kershner 1936; Zemlyakov and Katok 1975; Boldrighini et al. 1978); in special cases of rational billard tables, including the square and rectangle, this unfolding leads to a tiling of the plane and is amenable to illustration.
@@ 3 Properties of the square billiard
The alternative representation in Figure 4 is the basis for the classical proof of existence and classification of periodic orbits on the square billiard, which uses the notion of geodesic on $\mathbb{T}^{2}$, that is, a length-minimising curve. In the usual Euclidean geometry, geodesics are straight lines, and $\mathbb{T}^{2}$ inherits this geometry by way of the construction illustrated in Figure 4. The theory of geodesics tell us that all straight lines with rational slopes correspond to closed geodesics on $\mathbb{T}^{2}$; for example, see (Petersen 1983). Therefore, if a trajectory on the square billiard is periodic, then the slope must be rational.
Remark 2.
The theory of geodesics also states that all straight lines with irrational slopes correspond to geodesics that densely fill the torus. This dichotomy is known as Veech dichotomy in the literature; it implies that the square billiard is ergodically optimal (Veech 1989). We refer again to (Petersen 1983) for details.
The classical proof has the advantage that it extends readily to higher dimensions by using analogous arguments (Tabachnikov 2005). However, this approach ignores the possibility that a billiard trajectory with rational slope terminates at one of the four vertices; indeed, the condition that the slope be rational is not sufficient to guarantee periodicity. In this paper, we analyse the trichotomy that includes the possibility of singular orbits for a square billiard. We use the technique of unfolding in $\mathbb{R}^{2}$ as an alternative and more intuitive way to prove important properties of the square billiard, such as the non-existence of periodic orbits with odd periods (Davis 2016) and what the period of a billiard trajectory will be for a given rational slope. We will also count and classify all different equivalence classes of periodic orbits that have a given period associated with a particular rational slope.
@@ 3.1 Periods of periodic orbits for the square billiard are even
Proof.
Suppose for the sake of contradiction that a period-three orbit exists. To focus the mind, Figure 5 illustrates this hypothetical billiard trajectory on the square ABCD. The three different points of collision are denoted $P_{0}$, $P_{1}$ and $P_{2}$, and their associated angles are $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$, respectively; here, angles with the same label are supposed to be equal, even though the image may suggest otherwise. Note that any period-three orbit must collide with three different sides of the square, because it is impossible for a billiard trajectory to incur two consecutive collisions on the same side. Since we can always rotate the table and relabel the vertices, the illustration in Figure 5 is representative for any period-three orbit. In other words, without loss of generality, we may assume that $P_{0}$ lies on the side AB, and the other two, successive points of collision $P_{1}$ and $P_{2}$ lie on the sides BC and CD, respectively. Recall from Section 2 that the billiard trajectory is uniquely generated by the pair $\langle P_{0},\alpha_{0}\rangle$, and Proposition 2.1 implies that we must have $\alpha_{1}=\frac{\pi}{2}-\alpha_{0}$ and $\alpha_{2}=\alpha_{0}$. Furthermore, the sum of the angles in the quadrilateral formed by $P_{0}$, $P_{2}$, and the two vertices D and A of the square table must be $2\pi$, which implies that $\alpha_{0}+\alpha_{2}=\pi$. Therefore, if a period-three orbit exists then
[…]
However, the pair $\langle P_{0},\frac{\pi}{2}\rangle$ generates a period-two orbit for all $P_{0}\in(0,1)$, because the billiard trajectory will bounce back and forth between the two points $P_{0}$ and $P_{2}$; a contradiction. ∎
@@
Proof.
Consider a periodic orbit for the square ABCD generated by a pair $\langle P_{0},\alpha_{0}\rangle$ with $P_{0}\in[0,1)$ and $\alpha\in(0,\frac{\pi}{2}]$. Note that $P_{0}\neq 0$, because a periodic billiard trajectory that starts at a vertex must also return to this vertex, which would make it a singular orbit. The unfolding on $\mathbb{R}^{2}$ of such a periodic orbit is a straight line with slope $\tan{(\alpha_{0})}$ that passes through the point at distance $P_{0}$ from the vertex A at the origin, and such that it intersects the tiling of $\mathbb{R}^{2}$ along other (although not necessarily all) sides AB at distances $P_{0}$ from the shifted vertex A in translated copies of the original table ABCD.
@@
Furthermore, Definition 3.1 states that a translated copy of the billiard table with orientation ABCD is only obtained when the (integer) coordinates of the shifted vertex A are both even, that is, $p$ and $q$ must both be even. The period of the periodic orbit is given by the total numer of distinct collisions; we claim that this number is $p+q$, and since $p$ and $q$ are even, the period of the periodic orbit is even. It remains to prove that the total numer, say $K$, of distinct collisions is $p+q$. Recall that the technique of unfolding transforms the billiard trajectory into a straight line by reflecting the table at points of collision, instead of changing the direction of motion. Hence, $K$ is at most equal to the number of reflections, that is, $K\leq p+q$. Now suppose $K<p+q$, which means that the billiard trajectory is repeated after $K$ collisions. However, $p$ and $q$ are, by definition, the number of horizontal and vertical reflections needed to pass through a translated copy of $P_{0}$ with the correct orientation for the first time. Hence, $P_{\ell}\neq P_{0}$ for all $0<\ell<p+q$; therefore, $K=p+q$ and the period is even. ∎
@@ 3.2 Construction of periodic orbits for the square billiard
If the trajectory generated by a pair $\langle P_{0},\alpha_{0}\rangle$ is periodic, then the unfolded trajectory will intersect shifted copies of the side AB infinitely many times at distances $P_{0}$. The (rational) slope $\tan{(\alpha_{0})}$ of this trajectory can be determined using any such shifted copy: if a periodic orbit is obtained after $p$ horizontal and $q$ vertical reflections then the same periodic orbit is obtained after $d\,p$ horizontal and $d\,q$ vertical reflections, for any $d\in\mathbb{N}$, because $\tan{(\alpha_{0})}=\frac{p}{q}=\frac{d\,p}{d\,q}$. For the same reason, if $d\in\{3,\ldots,\min{(p,q)}\}$ divides both $p$ and $q$ then there exists an intersection at distance $P_{0}$ on a shifted copy of the side AB that lies closer to the original AB. Of course, $d=2$ divides $p$ and $q$, because they are both even, but the corresponding point $(P_{0}+\frac{p}{2},\,\frac{q}{2})$ does not necessarily lie on the side AB, or if it does, this side may have the reflected orientation BA instead. Inspired by Baxter and Umble (2008) who give a similar definition for the equilateral triangle, we formalise the notion of least period of a periodic unfolded trajectory.
Definition 3.4.
@@
For sufficiently large $K$, we expect there to be other pairs of co-prime numbers that sum to $\frac{K}{2}$, that is, we expect there to be more equivalence classes and different types of periodic orbits with the same period $K\in 2\,\mathbb{N}$. To determine how many possible decompositions there are, we use Euler’s totient function from number theory (Loomis et al. 2008). For any $N\in\mathbb{N}$, Euler’s totient function $\varphi(N)$ counts the number of natural numbers $n\in\{1,\ldots,N\}$ such that $\gcd{(n,N)}=1$. The integers $n$ that satisfy this property are referred to as totatives of $N$.
Proposition 3.5.
@@
Proof.
Consider $m,n\in\mathbb{N}$ with $\gcd{(m,n)}=1$ and define $N=m+n$. By definition, $\gcd{(m,n)}=\gcd{(m,\,m+n)}=\gcd{(m,N)}$. However, $\gcd{(m,n)}=1$, so $\gcd{(m,N)}=1$, which means that $m$ is a totative of $N$. Therefore, each ordered pair $(m,n)$ is determined by whether $m$ is a totative of $N$, and the total number of possibilities is $\varphi(N)$. ∎
@@
[…]
can be derived using the discrete Fourier transform (Schramm 2008); unfortunately, finding the greatest common divisors $\gcd{(k,N)}$ is also computationally as complex as finding prime factors.

### v2

@@ Abstract
Keywords: periodic orbit; mathematical billiard; Euler’s totient function.
@@ 1 Introduction
Mathematical billiards shares many similarities to the game of billiards in reality. Fundamentally, both comprise a billiard ball moving in a straight line until it strikes a side of the table. Mathematical billiards ignores friction, takes the billiard ball as a point mass and assumes purely elastic collisions. As a consequence, the billiard trajectory will satisfy the law of reflection (the angle of incidence equals the angle of reflection) when it collides with a boundary. Therefore, any billiard trajectory is uniquely determined by its initial position and direction of motion. A real billiard table is almost always rectangular, but mathematical billiard tables can have arbitrary shapes and dimensions. First posed as Alhazen’s problem in optics by Ptolemy in 150 AD [3], mathematical billiards is much more than a fun and interesting game; it turns out that the mere shape of the table distinguishes mathematical billiards into three different classes—elliptic, hyperbolic, and parabolic—which are well known as classes in different fields of mathematics, not the least in dynamical systems theory [25, 33]. Mathematical billiards has extensively been studied from the perspective of ergodic theory in dynamical systems, as well as algebraic geometry (moduli spaces) and Teichmüller theory; for example, see the textbooks by Tabachnikov [39], Chernov and Markarian [10], or Rozikov [32]. Nevertheless, a myriad of unsolved open problems in mathematical billiards remain [5, 13, 19, 22]. Indeed, there are many fundamental theory questions in mathematics that can be recast as a billiard problem [34, 35], as well as a plethora of applications, including but not limited to the computation of $\pi$ [18], mechanics [1], quantum computing [17], pouring problems [30], Benford’s Law [39], diffusion in the Lorentz Gas [14] and the Riemann hypothesis [8].
@@
Obviously, singular orbits only exist on tables with vertices and they are often ignored or classified as non-periodic [7]. If a billiard trajectory is periodic, its period is given by the total number of unique collisions. A natural question to ask is whether periodic orbits always exist or whether this depends on the shape of the billiard table. Furthermore, what periods are possible? In fact, the very question whether periodic orbits exist on any polygon is still an open problem! It is listed as Problem 3 in Katok’s five most resistant problems in dynamics [24]. Even the case of a triangular billiard table is considered impenetrable, although the Triangular Billiards Conjecture, which states that every triangular table has a periodic billiard path, is widely believed to be true [38].
The question about which periods are possible has been treated for specific polygonal billiard tables. In particular, Baxter and Umble [4] classified all possible periodic orbits for an equilateral triangle. In this paper, we consider square or rectangular billiard tables and we similarly give a complete classification of which periods and types of periodic orbits are possible for a square or rectangle. The square is perhaps the simplest piecewise-smooth billiard table, which means that it is introduced early on in billiard textbooks, typically alongside the smooth, circular billiard; for example, see [10, 32, 39]. As a consequence, results for the square billiard tend to be skimmed over, despite its behaviour being distinctly different from circular and other smooth, convex tables. Existence of periodic orbits on the square and rectangular billiard is easily verified by explicit examples. Figure 1 shows three examples of periodic orbits on the square. Here, the square table is represented by the (blue) boundary with vertices A, B, C and D, the periodic orbits are indicated as directed (black) lines that collide with the boundary at points identified by thick (red) dots. By counting these points of collision, we see that the periods are two, four and six in the respective panels (a), (b) and (c). Note that the period-two orbit in panel (a) is readily produced on a rectangle as well; for the period-four and -six orbits this is less obvious.
@@
Interestingly, none of the textbooks show that the square and rectangular billiard tables do not admit periodic orbits with odd periods. This can be proven using cutting sequences, as done by Davis [11], but we provide a different proof that harmonises well with the periodicity classification presented in this paper. Furthermore, we also discuss how to construct periodic orbits with prespecified periods and connect our classification of periodic orbits with Euler’s totient function from number theory [27].
This paper is organised as follows. In the next section, we define the different classes of periodic orbits and clarify what we mean when two periodic orbits with the same period are different. Here, we also explain the powerful technique of unfolding [16, 41] that is used in the classical proof of the existence of periodic orbits on the square or rectangular billiard table. We present alternative and intuitive proofs for periodicity and non-existence of odd periodic orbits in Section 3. In Section 3.2, we count and fully classify all possible periodic orbits. We treat singular orbits in Section 3.3, where we show how different types of singular orbits form boundaries between different families of periodic orbits. In Section 4, we extend the results for the square to the rectangular billiard table. We conclude in Section 5 with a discussion of future work.
@@ 2 Setting and definitions
Proof. Since $\alpha_{0}\in(0,\frac{\pi}{2}]$, the outgoing line from $P_{0}$ on the side AB will either end on the adjacent side BC or on the opposite side CD. The sides AB and CD are parallel, because ABCD is either a square or a rectangle. Hence, if the outgoing line from $P_{0}$ ends on CD, is does so at the same angle $\alpha_{0}$; see Figure 1(a) and 1(c) for examples. If the outgoing line from $P_{0}$ ends on the adjacent side BC then it does so at an angle $\alpha_{1}$ such that $\alpha_{0}+\frac{\pi}{2}+\alpha_{1}=\pi$, because the outgoing line forms a right triangle with the sides AB and BC; see Figure 1(b) and 1(c) for examples. Therefore, we have $\alpha_{1}=\frac{\pi}{2}-\alpha_{0}$ as claimed. The end point of the outgoing line from $P_{0}$ is another collision point $P_{1}$, unless it is a vertex, in which case the billiard trajectory terminates and the result holds. By rotating the table and relabelling the vertices, we can consider the billiard trajectory as being generated by the pair $\langle P_{1},\alpha_{1}\rangle$, with $\alpha_{1}=\alpha_{0}$ or $\alpha_{1}=\frac{\pi}{2}-\alpha_{0}$. Using the same arguments, we find that a collision with the (original) side DA also occurs at angle $\frac{\pi}{2}-\alpha_{0}$. It follows that all collisions with sides AB and CD occur at angle $\alpha_{0}$ and all collisions with sides BC and DA occur at angle $\frac{\pi}{2}-\alpha_{0}$.
$\square$
@@
Definition 2.2 (Equivalence class $\mathcal{C}_{K}(p)$ of period-$K$ orbits)
@@ 2.1 Unfolding of billiard trajectories
Following a billiard trajectory on a table ABCD can be difficult if there are many collisions, because of a myriad of intersecting lines. The technique of unfolding transforms the billiard trajectory: rather than reflecting collisions with the sides, the entire table reflects, so that the billiard trajectory remains a straight line. For so-called rational, planar, polygonal billiard tables, which have polygonal angles that are rational multiples of $\pi$, this approach relates mathematical billards to the theory of geodesic flow on Riemann surfaces [16, 41, 6]; in special cases of rational billard tables, including the square and rectangle, this unfolding leads to a tiling of the plane and is amenable to illustration.
@@ 3 Properties of the square billiard
The alternative representation in Figure 4 is the basis for the classical proof of existence and classification of periodic orbits on the square billiard, which uses the notion of geodesic on $\mathbb{T}^{2}$, that is, a length-minimising curve. In the usual Euclidean geometry, geodesics are straight lines, and $\mathbb{T}^{2}$ inherits this geometry by way of the construction illustrated in Figure 4. The theory of geodesics tell us that all straight lines with rational slopes correspond to closed geodesics on $\mathbb{T}^{2}$; for example, see [29]. Therefore, if a trajectory on the square billiard is periodic, then the slope must be rational.
Remark 2
The theory of geodesics also states that all straight lines with irrational slopes correspond to geodesics that densely fill the torus. This dichotomy is known as Veech dichotomy in the literature; it implies that the square billiard is ergodically optimal [40]. We refer again to [29] for details.
The classical proof has the advantage that it extends readily to higher dimensions by using analogous arguments [39]. However, this approach ignores the possibility that a billiard trajectory with rational slope terminates at one of the four vertices; indeed, the condition that the slope be rational is not sufficient to guarantee periodicity. In this paper, we analyse the trichotomy that includes the possibility of singular orbits for a square billiard. We use the technique of unfolding in $\mathbb{R}^{2}$ as an alternative and more intuitive way to prove important properties of the square billiard, such as the non-existence of periodic orbits with odd periods [11] and what the period of a billiard trajectory will be for a given rational slope. We will also count and classify all different equivalence classes of periodic orbits that have a given period associated with a particular rational slope.
@@ 3.1 Periods of periodic orbits for the square billiard are even
Proof. Suppose for the sake of contradiction that a period-three orbit exists. To focus the mind, Figure 5 illustrates this hypothetical billiard trajectory on the square ABCD. The three different points of collision are denoted $P_{0}$, $P_{1}$ and $P_{2}$, and their associated angles are $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$, respectively; here, angles with the same label are supposed to be equal, even though the image may suggest otherwise. Note that any period-three orbit must collide with three different sides of the square, because it is impossible for a billiard trajectory to incur two consecutive collisions on the same side. Since we can always rotate the table and relabel the vertices, the illustration in Figure 5 is representative for any period-three orbit. In other words, without loss of generality, we may assume that $P_{0}$ lies on the side AB, and the other two, successive points of collision $P_{1}$ and $P_{2}$ lie on the sides BC and CD, respectively. Recall from Section 2 that the billiard trajectory is uniquely generated by the pair $\langle P_{0},\alpha_{0}\rangle$, and Proposition 2.1 implies that we must have $\alpha_{1}=\frac{\pi}{2}-\alpha_{0}$ and $\alpha_{2}=\alpha_{0}$. Furthermore, the sum of the angles in the quadrilateral formed by $P_{0}$, $P_{2}$, and the two vertices D and A of the square table must be $2\pi$, which implies that $\alpha_{0}+\alpha_{2}=\pi$. Therefore, if a period-three orbit exists then
$\left.\begin{array}[]{rcrcl}\alpha_{0}&+&\alpha_{2}&=&\pi\\ &&\alpha_{2}&=&\alpha_{0}\end{array}\right\}\implies 2\,\alpha_{0}=\pi\Longleftrightarrow\alpha_{0}=\tfrac{\pi}{2}.$
However, the pair $\langle P_{0},\frac{\pi}{2}\rangle$ generates a period-two orbit for all $P_{0}\in(0,1)$, because the billiard trajectory will bounce back and forth between the two points $P_{0}$ and $P_{2}$; a contradiction.
$\square$
@@
Proof. Consider a periodic orbit for the square ABCD generated by a pair $\langle P_{0},\alpha_{0}\rangle$ with $P_{0}\in[0,1)$ and $\alpha\in(0,\frac{\pi}{2}]$. Note that $P_{0}\neq 0$, because a periodic billiard trajectory that starts at a vertex must also return to this vertex, which would make it a singular orbit. The unfolding on $\mathbb{R}^{2}$ of such a periodic orbit is a straight line with slope $\tan{(\alpha_{0})}$ that passes through the point at distance $P_{0}$ from the vertex A at the origin, and such that it intersects the tiling of $\mathbb{R}^{2}$ along other (although not necessarily all) sides AB at distances $P_{0}$ from the shifted vertex A in translated copies of the original table ABCD.
@@
Furthermore, Definition 3.1 states that a translated copy of the billiard table with orientation ABCD is only obtained when the (integer) coordinates of the shifted vertex A are both even, that is, $p$ and $q$ must both be even. The period of the periodic orbit is given by the total numer of distinct collisions; we claim that this number is $p+q$, and since $p$ and $q$ are even, the period of the periodic orbit is even. It remains to prove that the total numer, say $K$, of distinct collisions is $p+q$. Recall that the technique of unfolding transforms the billiard trajectory into a straight line by reflecting the table at points of collision, instead of changing the direction of motion. Hence, $K$ is at most equal to the number of reflections, that is, $K\leq p+q$. Now suppose $K<p+q$, which means that the billiard trajectory is repeated after $K$ collisions. However, $p$ and $q$ are, by definition, the number of horizontal and vertical reflections needed to pass through a translated copy of $P_{0}$ with the correct orientation for the first time. Hence, $P_{\ell}\neq P_{0}$ for all $0<\ell<p+q$; therefore, $K=p+q$ and the period is even.
$\square$
@@ 3.2 Construction of periodic orbits for the square billiard
If the trajectory generated by a pair $\langle P_{0},\alpha_{0}\rangle$ is periodic, then the unfolded trajectory will intersect shifted copies of the side AB infinitely many times at distances $P_{0}$. The (rational) slope $\tan{(\alpha_{0})}$ of this trajectory can be determined using any such shifted copy: if a periodic orbit is obtained after $p$ horizontal and $q$ vertical reflections then the same periodic orbit is obtained after $d\,p$ horizontal and $d\,q$ vertical reflections, for any $d\in\mathbb{N}$, because $\tan{(\alpha_{0})}=\frac{p}{q}=\frac{d\,p}{d\,q}$. For the same reason, if $d\in\{3,\ldots,\min{(p,q)}\}$ divides both $p$ and $q$ then there exists an intersection at distance $P_{0}$ on a shifted copy of the side AB that lies closer to the original AB. Of course, $d=2$ divides $p$ and $q$, because they are both even, but the corresponding point $(P_{0}+\frac{p}{2},\,\frac{q}{2})$ does not necessarily lie on the side AB, or if it does, this side may have the reflected orientation BA instead. Inspired by Baxter and Umble [4] who give a similar definition for the equilateral triangle, we formalise the notion of least period of a periodic unfolded trajectory.
Definition 3.4
@@
For sufficiently large $K$, we expect there to be other pairs of co-prime numbers that sum to $\frac{K}{2}$, that is, we expect there to be more equivalence classes and different types of periodic orbits with the same period $K\in 2\,\mathbb{N}$. To determine how many possible decompositions there are, we use Euler’s totient function from number theory [27]. For any $N\in\mathbb{N}$, Euler’s totient function $\varphi(N)$ counts the number of natural numbers $n\in\{1,\ldots,N\}$ such that $\gcd{(n,N)}=1$. The integers $n$ that satisfy this property are referred to as totatives of $N$.
Proposition 3.5
@@
Proof. Consider $m,n\in\mathbb{N}$ with $\gcd{(m,n)}=1$ and define $N=m+n$. By definition, $\gcd{(m,n)}=\gcd{(m,\,m+n)}=\gcd{(m,N)}$. However, $\gcd{(m,n)}=1$, so $\gcd{(m,N)}=1$, which means that $m$ is a totative of $N$. Therefore, each ordered pair $(m,n)$ is determined by whether $m$ is a totative of $N$, and the total number of possibilities is $\varphi(N)$.
$\square$
@@
$\varphi(N)=\sum_{k=1}^{N}\gcd{(k,N)}\cos{\left(2\pi\tfrac{k}{N}\right)},$
can be derived using the discrete Fourier transform [36]; unfortunately, finding the greatest common divisors $\gcd{(k,N)}$ is also computationally as complex as finding prime factors.